OpenStudy (solomonzelman):

Here I have some latex, but not naturally produced. (I will elaborate) I am using tricks, but maybe there is that function/latex designed for it...

OpenStudy (solomonzelman):

I will label in red what notations have I used a trick for, but would like to see if there is a natural way for it. \(\large\color{slate}{\displaystyle\int\limits_{a}^{b}f(x)~dx=~F(x)~\color{red}{{{\Huge |}_{a}^{b}}}}\) ``` \(\large\color{slate}{\displaystyle\int\limits_{a}^{b}f(x)~dx=~F(x)~ \color{green}{{ {\Huge |}_{a}^{b} }}}\) ``` this, \(\large\color{green}{ \Uparrow }\) (this is what I need to know) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ Another suggestion from poster (me). 1.5 spacing like in a word doc: `\(\LARGE\color{white}{ \left| \right| }\)` using something white to widen between-the-line spaces. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ then I need another 2 things.